# A quantum information method for early universe with non-trivial sound speed

### v1

@@ Abstract
The growing application of two-mode squeezed states in cosmology underscores the increasing relevance of quantum information concepts across various domains. Among these, Krylov complexity, a measure of operator growth in quantum dynamical systems based on the Lanczos algorithm, has garnered significant attention Parker:2018yvk . It has been widely applied in areas such as the SYK model in high-energy physics Rabinovici:2020ryf ; Jian:2020qpp ; He:2022ryk , the Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , conformal field theory Dymarsky:2021bjq ; Caputa:2021ori ; Kundu:2023hbk , quantum field theory He:2024hkw ; He:2024xjp ; He:2025guu , and topological phases Caputa:2022eye . In generic settings, Krylov complexity exhibits exponential growth, even in integrable systems, as verified in Camargo:2022rnt ; He:2024xjp . Moreover, Krylov complexity has been extended to open quantum systems Bhattacharyya:2025lsc ; Zhai:2024tkz ; Carolan:2024wov ; Bhattacharya:2023zqt ; Liu:2022god ; Bhattacharjee:2022lzy . It also plays a role in the holographic dictionary Li:2025fqz ; Aguilar-Gutierrez:2025kmw ; Heller:2024ldz ; Das:2024tnw ; Rabinovici:2023yex , with notable connections to JT gravity Bhattacharyya:2025gvd ; Ambrosini:2024sre and dilaton physics Chattopadhyay:2023fob . For the most recent developments in this rapidly evolving topic, we refer readers to Refs. Rabinovici:2025otw ; Nandy:2024evd .
In this work, we will employ the framework of Krylov complexity to study the early universe, encompassing inflation, the radiation-dominated (RD), and matter-dominated (MD) eras, within models featuring a non-trivial sound speed. While several previous studies have explored computational complexity during inflation Choudhury:2020hil ; Bhargava:2020fhl ; Lehners:2020pem ; Bhattacharyya:2020rpy ; Adhikari:2021ked , applications of Krylov complexity to the broader evolutionary history of the early universe remain scarce. The work of Adhikari:2022oxr represents the first attempt to apply Krylov complexity to inflation, though it relies on a closed-system approach. In reality, however, the early universe constitutes an open quantum system. Recently, Li:2024kfm adopted an open-system approach to construct the wave function of the inflationary universe, revealing its strongly dissipative nature. Building on this perspective, the present study takes the wave function in open systems and the generalized Lanczos algorithm as its foundational starting point.
@@ II.2 Non-trivial sound speed during different epochs
$\begin{cases}\mbox{Inflation:}&a(\eta)=-\tau^{-1}\\ &c_{S}^{2}=1-2\xi[1-\cos{(k10^{-y})}]\\ \mbox{RD:}&a(\eta)=\eta\\ &c_{S}^{2}=1-2\xi[1-\cos{(k10^{y})}]\\ \mbox{MD:}&a(\eta)=\eta^{2}\\ &c_{S}^{2}=1-2\xi[1-\cos{(k10^{\frac{y}{2}})}]\end{cases}$ (11)
@@ V.1 The wave function for the open system
Here, we could see that the OTMSS is more genric and more complete compared to two-mode squeezed state, which incorporates the information of open system part for the Hamiltonian since its information is encoded in $u_{2}$. And the derivation of OTMSS is only related to the second kind Meixner polynomial Hetyei_2009 , whose formula is
@@ V.2 Evolution of $r_{k}$ and $\phi_{k}$
$\begin{split}\frac{H_{0}}{\ln 10}10^{-\frac{y}{2}(1+3\omega)}\frac{dr_{k}}{dy}=&\frac{|1-u_{1}^{2}|^{\frac{1}{2}}\sinh{2r_{k}}(\frac{10^{2y}V_{\phi\phi}}{2k}-\frac{k}{2}+\frac{kc_{S}^{2}}{2})\sin{2\phi_{k}}}{\sin{2r_{k}}+2u_{2}\cosh^{2}{r_{k}}}\\ &-\frac{|1-u_{1}^{2}|^{\frac{1}{2}}\sinh{2r_{k}}(H_{0}10^{-\frac{y}{2}(1+3\omega)}-\frac{\xi k\sin{k\tau}}{1-2\xi[1-\cos{k\tau}]})\cos{2\phi_{k}}}{\sin{2r_{k}}+2u_{2}\cosh^{2}{r_{k}}}\\ &-\frac{u^{\prime}_{2}\sinh{2r_{k}}}{\sin{2r_{k}}+2u_{2}\cosh^{2}{r_{k}}}\end{split}$ (47)
$\begin{split}\frac{H_{0}}{\ln 10}10^{-\frac{y}{2}(1+3\omega)}\frac{d\phi_{k}}{dy}=&-\frac{1}{2}(k+kc^{2}_{S}+\frac{10^{2y}V_{\phi\phi}}{k})+\frac{1}{2}[(\frac{a^{2}V_{\phi\phi}}{2k}-\frac{k}{2}+\frac{kc^{2}_{S}}{2})\cos{2\phi_{k}}\\ &+(H_{0}10^{-\frac{y}{2}(1+3\omega)}-\frac{\xi k\sin{k\tau}}{1-2\xi[1-\cos{k\tau}]})\sin{2\phi_{k}}]\\ &(|1-u_{1}^{2}|^{-\frac{1}{2}}(u_{2}+\coth{r_{k}})+|1-u_{1}^{2}|^{\frac{1}{2}}\frac{\tanh{r_{k}}}{1+u_{2}\tanh{r_{k}}})\end{split}$ (48)
@@
$\begin{split}\begin{cases}&u_{2}=\frac{k}{2}+\frac{kc_{S}^{2}}{2}+\frac{10^{2y}V_{,\phi\phi}}{2k},\\ &u_{2}^{\prime}=H_{0}10^{-\frac{y}{2}(1+3\omega)}\frac{10^{2y}V_{,\phi\phi}}{k}+\frac{10^{2y}V_{,\phi\phi}^{\prime}}{2k}+kc_{S}c_{S}^{\prime}\\ &\ |1-u_{1}^{2}|=(\frac{10^{2y}V_{,\phi\phi}}{2k}-\frac{k}{2}+\frac{kc_{S}^{2}}{2})^{2}+(H_{0}10^{-\frac{y}{2}(1+3\omega)}-\frac{\xi k\sin{k\tau}}{1-2\xi[1-\cos{k\tau}]})^{2}.\end{cases}\end{split}$
@@
Inflation: The overall trend of $r_{k}$ is exponential growth, consistent with previous findings for the case $\xi=0$, which corresponds to single-field inflation Zhai:2024odw . Increasing $\xi$ enhances the amplitude of $r_{k}$. During the RD era, $r_{k}$ exhibits similar behavior across different values of $\xi$: it grows and eventually saturates at a constant value. Higher values of $\xi$ further amplify the saturation amplitude of $r_{k}$. In the MD era, the amplitude of $r_{k}$ is also enhanced with larger $\xi$. The overall behavior remains consistent: $r_{k}$ reaches a nearly constant value before resuming growth for $y>4.5$.
@@ V.3 Evolution of Krylov complexity and Krylov entropy
Krylov entropy quantifies the disorder or spread of an operator’s evolution within the Krylov space He:2024xjp , measuring the ”disorder” of the Lanczos coefficients. This paper demonstrates its utility in distinguishing between standard single-field inflation and models with non-trivial sound speed. Our analysis reveals that Krylov entropy acts as a powerful diagnostic tool to distinguish between standard single-field inflation and models with a non-trivial sound speed ($c_{S}\neq 1$). A key finding is the significant impact of the parameter $\phi_{k}$, which sets the scale of the feature in the sound speed—on the Krylov entropy. For instance, in inflationary scenarios, a value of $\xi=0.02$ (where $\xi$ is a parameter derived from $\phi_{k}$) emerges as a critical threshold, producing a stark contrast in the entropy profile. This distinction is not limited to inflation; it is also evident in the subsequent RD and MD eras. The behavior of the Krylov entropy in these epochs further differentiates the standard cosmological model from one with a non-trivial sound speed. The sensitivity of Krylov entropy to the feature scale $\phi_{k}$ establishes it as a robust criterion for discriminating between standard single-field inflation and models with non-trivial kinetic terms characterized by $c_{S}$.
@@ VI Summary and discussion
Fig. 7 clearly shows the evolution of Krylov complexity: it grows exponentially during inflation, decays toward a constant in the radiation-dominated (RD) era, and exhibits a peak in the matter-dominated (MD) epoch. While all cases follow a similar trend, a key difference from Parker:2018yvk is that, although the early universe remains a chaotic system, the Krylov complexity does not saturate to a fixed value, both for non-trivial sound speed and standard cases, consistent with our previous findings Li:2024kfm .

### v2

@@ Abstract
The growing application of two-mode squeezed states in cosmology underscores the increasing relevance of quantum information concepts across various domains. Among these, Krylov complexity, a measure of operator growth in quantum dynamical systems based on the Lanczos algorithm, has garnered significant attention Parker:2018yvk . It has been widely applied in areas such as the SYK model in high-energy physics Rabinovici:2020ryf ; Jian:2020qpp ; He:2022ryk , the Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , conformal field theory Dymarsky:2021bjq ; Caputa:2021ori ; Kundu:2023hbk , quantum field theory He:2024hkw ; He:2024xjp ; He:2025guu , and topological phases Caputa:2022eye . In generic settings, Krylov complexity exhibits exponential growth, even in integrable systems, as verified in Camargo:2022rnt ; He:2024xjp . Moreover, Krylov complexity has been extended to open quantum systems Bhattacharyya:2025lsc ; Zhai:2024tkz ; Carolan:2024wov ; Bhattacharya:2023zqt ; Liu:2022god ; Bhattacharjee:2022lzy . It also plays a role in the holographic dictionary Li:2025fqz ; Aguilar-Gutierrez:2025kmw ; Heller:2024ldz ; Das:2024tnw ; Rabinovici:2023yex , with notable connections to JT gravity Bhattacharyya:2025gvd ; Ambrosini:2024sre and dilaton physics Chattopadhyay:2023fob . For the most recent developments in this rapidly evolving topic, we refer readers to Refs. Rabinovici:2025otw ; Nandy:2024evd . Essentially speaking, the study of complexity in the early universe is fundamentally an investigation into quantum information within a non-inertial framework. Because the expansion of spacetime renders observers non-inertial relative to distant cosmological scales, it is crucial to consider the perspective of quantum information in non-inertial systems. This approach aligns with recent developments in relativistic quantum information, such as those discussed in Wu:2022xwy ; Wu:2023spa ; Wu:2023sye ; Li:2025bzd ; Li:2025jlu ; Liu:2024wpa ; Liu:2025hcx .
In this work, we will employ the framework of Krylov complexity to study the early universe, encompassing inflation, the radiation-dominated (RD), and matter-dominated (MD) eras, within models featuring a non-trivial sound speed. While several previous studies have explored computational complexity during inflation Choudhury:2020hil ; Bhargava:2020fhl ; Lehners:2020pem ; Bhattacharyya:2020rpy ; Adhikari:2021ked . The work of Adhikari:2022oxr represents the first attempt to apply Krylov complexity to inflation, though it relies on a closed-system approach. In reality, however, the early universe constitutes an open quantum system. Recently, Li:2024kfm adopted an open-system approach to construct the wave function of the inflationary universe, revealing its strongly dissipative nature. Building on this perspective, the present study takes the wave function in open systems and the generalized Lanczos algorithm as its foundational starting point.
@@ II.2 Non-trivial sound speed during different epochs
$\begin{cases}\mbox{Inflation:}&a(\eta)=-\eta^{-1}\\ &c_{S}^{2}=1-2\xi[1-\cos{(k10^{-y})}]\\ \mbox{RD:}&a(\eta)=\eta\\ &c_{S}^{2}=1-2\xi[1-\cos{(k10^{y})}]\\ \mbox{MD:}&a(\eta)=\eta^{2}\\ &c_{S}^{2}=1-2\xi[1-\cos{(k10^{\frac{y}{2}})}]\end{cases}$ (11)
@@ V.1 The wave function for the open system
During the RD and MD eras, the dissipation coefficient $u_{2}$ remains significantly small ($u_{2}\ll 1$), placing the universe in a weak dissipative phase. Consequently, while the OTMSS captures essential non-unitary information, the results in these later epochs do not deviate drastically from the ’closed system’ approximation, as the open-system corrections remain perturbative. Here, we could see that the OTMSS is more genric and more complete compared to two-mode squeezed state, which incorporates the information of open system part for the Hamiltonian since its information is encoded in $u_{2}$. And the derivation of OTMSS is only related to the second kind Meixner polynomial Hetyei_2009 , whose formula is
@@ V.2 Evolution of $r_{k}$ and $\phi_{k}$
$\begin{split}\frac{H_{0}}{\ln 10}10^{-\frac{y}{2}(1+3\omega)}\frac{dr_{k}}{dy}=&\frac{|1-u_{1}^{2}|^{\frac{1}{2}}\sinh{2r_{k}}(\frac{10^{2y}V_{\phi\phi}}{2k}-\frac{k}{2}+\frac{kc_{S}^{2}}{2})\sin{2\phi_{k}}}{\sin{2r_{k}}+2u_{2}\cosh^{2}{r_{k}}}\\ &-\frac{|1-u_{1}^{2}|^{\frac{1}{2}}\sinh{2r_{k}}(H_{0}10^{-\frac{y}{2}(1+3\omega)}-\frac{\xi k\sin{k\eta}}{1-2\xi[1-\cos{k\eta}]})\cos{2\phi_{k}}}{\sin{2r_{k}}+2u_{2}\cosh^{2}{r_{k}}}\\ &-\frac{u^{\prime}_{2}\sinh{2r_{k}}}{\sin{2r_{k}}+2u_{2}\cosh^{2}{r_{k}}}\end{split}$ (47)
$\begin{split}\frac{H_{0}}{\ln 10}10^{-\frac{y}{2}(1+3\omega)}\frac{d\phi_{k}}{dy}=&-\frac{1}{2}(k+kc^{2}_{S}+\frac{10^{2y}V_{\phi\phi}}{k})+\frac{1}{2}[(\frac{a^{2}V_{\phi\phi}}{2k}-\frac{k}{2}+\frac{kc^{2}_{S}}{2})\cos{2\phi_{k}}\\ &+(H_{0}10^{-\frac{y}{2}(1+3\omega)}-\frac{\xi k\sin{k\eta}}{1-2\xi[1-\cos{k\eta}]})\sin{2\phi_{k}}]\\ &(|1-u_{1}^{2}|^{-\frac{1}{2}}(u_{2}+\coth{r_{k}})+|1-u_{1}^{2}|^{\frac{1}{2}}\frac{\tanh{r_{k}}}{1+u_{2}\tanh{r_{k}}})\end{split}$ (48)
@@
$\begin{split}\begin{cases}&u_{2}=\frac{k}{2}+\frac{kc_{S}^{2}}{2}+\frac{10^{2y}V_{,\phi\phi}}{2k},\\ &u_{2}^{\prime}=H_{0}10^{-\frac{y}{2}(1+3\omega)}\frac{10^{2y}V_{,\phi\phi}}{k}+\frac{10^{2y}V_{,\phi\phi}^{\prime}}{2k}+kc_{S}c_{S}^{\prime}\\ &\ |1-u_{1}^{2}|=(\frac{10^{2y}V_{,\phi\phi}}{2k}-\frac{k}{2}+\frac{kc_{S}^{2}}{2})^{2}+(H_{0}10^{-\frac{y}{2}(1+3\omega)}-\frac{\xi k\sin{k\eta}}{1-2\xi[1-\cos{k\eta}]})^{2}.\end{cases}\end{split}$
@@
Inflation: The overall trend of $r_{k}$ is exponential growth, consistent with previous findings for the case $\xi=0$, which corresponds to single-field inflation Zhai:2024odw . Increasing $\xi$ enhances the amplitude of $r_{k}$. The enhancement of the squeezing amplitude $r_{k}$ in the presence of a non-trivial sound speed can be understood through the modified kinetic structure in Eq. (10). When $c_{S}^{2}$ deviates from unity, it effectively modulates the frequency of the Mukhannov-Sasaki variable. This modulation triggers a parametric resonance-like effect during the inflationary expansion, leading to the enhancement into the squeezing of quantum states compared to the standard canonical case ($c_{S}=1$). During the RD era, $r_{k}$ exhibits similar behavior across different values of $\xi$: it grows and eventually saturates at a constant value. Higher values of $\xi$ further amplify the saturation amplitude of $r_{k}$. In the MD era, the amplitude of $r_{k}$ is also enhanced with larger $\xi$. The overall behavior remains consistent: $r_{k}$ reaches a nearly constant value before resuming growth for $y>4.5$.
@@ V.3 Evolution of Krylov complexity and Krylov entropy
Krylov entropy quantifies the disorder or spread of an operator’s evolution within the Krylov space He:2024xjp , measuring the ”disorder” of the Lanczos coefficients. This paper demonstrates its utility in distinguishing between standard single-field inflation and models with non-trivial sound speed. Our analysis reveals that Krylov entropy acts as a powerful diagnostic tool to distinguish between standard single-field inflation and models with a non-trivial sound speed ($c_{S}\neq 1$). A key finding is the significant impact of the parameter $\phi_{k}$, which sets the scale of the feature in the sound speed—on the Krylov entropy. Numerical simulations reveal that $\xi=0.02$ acts as a critical threshold for the behavior of Krylov entropy. At this value, the oscillatory features introduced by the Sound Speed Resonance (SSR) become dominant enough to produce a distinct peak in entropy before horizon exit. While the specific numerical value may shift slightly with different choices of momentum $k$ (we have set $H_{0}=1$ for the entire early universe), the emergence of this ’peak-and-saturate’ behavior is a robust feature of the SSR model that distinguishes it from the monotonic growth seen in standard single-field inflation in inflationary period. This distinction is not limited to inflation; it is also evident in the subsequent RD and MD eras. The sensitivity of Krylov entropy to the feature scale $\phi_{k}$ establishes it as a robust criterion for discriminating between standard single-field inflation and models with non-trivial kinetic terms characterized by $c_{S}$.
@@ VI Summary and discussion
Fig. 7 clearly shows the evolution of Krylov complexity: it grows exponentially during inflation, decays toward a constant in the radiation-dominated (RD) era, and exhibits a peak in the matter-dominated (MD) epoch. While all cases follow a similar trend, a key difference from Parker:2018yvk is that, although the early universe remains a chaotic system, the Krylov complexity does not saturate to a fixed value, both for non-trivial sound speed and standard cases, consistent with our previous findings Li:2024kfm . Unlike many finite-dimensional chaotic systems where Krylov complexity eventually reaches a constant saturation value, our results show different trend. This distinction arises from the infinite-dimensional nature of the Hilbert space associated with the continuous mode evolution in an expanding universe.
@@ Acknowledgements
Data availability Statement
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.